OpenStudy (purplemexican):

Find the sum of the given arithmetic series. 24 + 48 + 72 + 96 + ... + 480 a. 5280 c. 10080 b. 9576 d. 5040

OpenStudy (anonymous):

\[480 = 24+ (n - 1) 24 \]

OpenStudy (anonymous):

\[S_n = \frac{ 19(24+480) }{ 2 } = ?\]

OpenStudy (purplemexican):

4788

OpenStudy (anonymous):

doesn't seem to be in your option

OpenStudy (purplemexican):

nope

OpenStudy (anonymous):

weird lol

OpenStudy (purplemexican):

try the maths again maybe it was a miscalculation

OpenStudy (anonymous):

true enough

OpenStudy (anonymous):

its \[S _{20} = \frac{ 20(24+480) }{ 2 } =5040\]

OpenStudy (purplemexican):

i think thats it

OpenStudy (purplemexican):

thank you

OpenStudy (anonymous):

yw